. If the numbers a, b, c, d , e form an A.P., then the value of a – 4b + 6c – 4d + e is …..

Grade:9

1 Answers

Arun
6 years ago
Dear Shivani
 

b = a +D

c = a +2D

d = a+3D

e = a+ 4D

where D is the common difference of the AP.

now put these values in the quest.

a - 4(a+D) +6(a+2D) -4(a+3D) + a+ 4D

=  0  ANS.

 

Regards

Arun (askIITians forum expert)